# ATLAS: Adaptive Self-Evolutionary Research Agent with Task-Distributed Multi-LLM Supporters

### v2

@@ A.3 Auxiliary Lemmas
$\hat{Z}(x_{t})=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta}(u_{t}(x_{t},y_{t})-\Delta u_{t}))$
$Z(x_{t})\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta}(\hat{u}_{t}(x_{t},y_{t})-u_{t}(x_{t},y_{t})))$
$\frac{\hat{Z}(x_{t})}{Z(x_{t})}=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{\hat{u}_{t}(x_{t},y_{t})-u_{t}(x_{t},y_{t})}{\beta}}$
Thus, the log term can be concluded as below
$\log{\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}}=\log\mathbb{E}_{y_{t}\sim\pi^{kl}(\cdot|x_{t})}[\exp{(-\frac{\Delta u_{t}}{\beta})}]+\frac{1}{\beta}(\Delta u_{t})$
Now we take the KL-divergence definition again
$D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))=\mathbb{E}_{y_{t}\sim\pi^{kl}}\biggl[\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}\biggr]=\log\mathbb{E}_{y_{t}\sim\pi^{kl}}\biggl[\exp{\bigl(-\frac{\Delta u_{t}}{\beta}\bigr)}\biggr]+\frac{1}{\beta}\mathbb{E}_{\pi^{kl}}[\Delta u_{t}]$
Let $\frac{\Delta u_{t}}{\beta}=X$. Then,
$D_{KL}=-\mathbb{E}[X]+\log\mathbb{E}_{y_{t}\sim\pi^{kl}}\biggl[e^{-(X-\mathbb{E}[X])}\biggr]+\mathbb{E}[X]$
Thus, $D_{KL}$ can be simplified as below
$D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))=\log\mathbb{E}_{y_{t}\sim\pi^{kl}}\bigl[e^{-(X-\mathbb{E}[X])}\bigr]$
Now, we derive the upper bound of $\Delta u_{t}$ with parameter estimate error. Since we defined $u_{t}$ in (3),
$\Delta u_{t}(y)=\theta^{\top}_{t}\phi(x_{t},y_{t})-\hat{\theta}^{\top}_{t}\phi(x_{t},y_{t})\leq\phi_{max}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}$
@@
$\frac{\Delta u_{t}}{\beta}\in\biggl[-\frac{\phi_{max}\|\theta_{t}-\hat{\theta}_{t}\|_{2}}{\beta},\frac{\phi_{max}\|\theta_{t}-\hat{\theta}_{t}\|_{2}}{\beta}\biggr]$
Let $a:=\frac{\phi_{max}\|\theta_{t}-\hat{\theta}_{t}\|_{2}}{\beta}$ so that $X\in[-a,a]$. Since $\mathbb{E}[X]\in[-a,a]$, the centered variable satisfies
@@
$D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))\leq\frac{\phi_{max}^{2}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}^{2}}{2\beta^{2}}$
A.4 Proof of regret analysis
We decompose the cumulative dynamic regret $R_{T}$ into Learning Error ($R^{error}_{T}$) and Reference Bias ($R^{bias}_{T}$).
Learning error
Now, we derive $J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})$.
$r_{t}=J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})=\sum_{y_{t}\in\mathcal{Y}}\pi^{kl}_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})-\sum_{y_{t}\in\mathcal{Y}}\pi_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})$
$\sum_{y_{t}\in\mathcal{Y}}\pi^{kl}_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})-\sum_{y_{t}\in\mathcal{Y}}\pi_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})\leq\sum_{y_{t}\in\mathcal{Y}}\biggl|\bigl(\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\bigr)\cdot u_{t}(x_{t},y_{t})\biggr|$
$\sum_{y_{t}\in\mathcal{Y}}\biggl|\bigl(\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\bigr)\cdot u_{t}(x_{t},y_{t})\biggr|\leq\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|\cdot|u_{t}(x_{t},y_{t})|$
@@ A.4 Proof of regret analysis
$\leq\|u_{t}(x_{t},\cdot)\|_{\infty}\sum_{y_{t}\in\mathcal{Y}}\Bigl|\pi_{t}^{kl}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\Bigr|$
$\text{norm inequality}\quad\|u\|_{\infty}\leq\|u\|_{2}\leq\|u\|_{1}$
We can write again as
$r_{t}\leq\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|\cdot\|u_{t}(x_{t},\cdot)\|_{\infty}\leq\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|\cdot\|u_{t}(x_{t},\cdot)\|_{2}$
We convert $\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|$ into L1 norm $\|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\|_{1}$ and using Assumption 1, derive the upper bound as below
$r_{t}\leq\theta_{max}\cdot\phi_{max}\cdot\|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\|_{1}$
Since the $\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})$ is measured in L1 norm, we convert the term into a KL-divergence term using Pinsker’s inequality and Lemma A.7,
$J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})\leq\sqrt{2D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))}\leq\frac{\theta_{max}\phi^{2}_{max}}{\beta}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}$
Summing over $t$ and applying Lemma A.6, together with $\sum^{T}_{t=1}V_{t,W}\leq W\;V_{T}$,
$R^{error}_{T}\leq\sum^{T}_{t=1}\frac{\theta_{max}\phi^{2}_{max}}{\beta}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}.$
From Lemma A.6, we finalize the cumulative learning error bound
$R^{error}_{T}\leq\frac{T\;\theta_{max}\phi^{2}_{max}}{\beta}\biggl[\frac{L_{\sigma}\phi_{\max}^{2}\,\,V_{t,W}}{m_{0}\cdot c}\;+\;\frac{\sqrt{\lambda+W\phi_{\max}^{2}}}{m_{0}cW}\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\;+\;\frac{\lambda\theta_{\max}}{m_{0}cW}\biggr]$
Since total variation is bounded by $V_{T}$ and $W=T^{\kappa}$, we write this upper bound with $\mathcal{O}$ function dependent on $T$ term, with probability at least $1-\delta$
@@
$R^{error}_{T}\leq\mathcal{O}\bigl(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log T}+T^{1-\kappa}).$
Reference Bias
We take the definition of $\pi^{kl}(y_{t}|x_{t})$ again
$\pi^{kl}_{t}(y_{t}|x_{t})=\frac{1}{Z(x)}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}u_{t}(x_{t},y_{t})}$
$\log\pi^{kl}_{t}(y_{t}|x_{t})=\log\pi_{ref,t}(y_{t}|x_{t})+\frac{1}{\beta}u_{t}(x_{t},y_{t})-\log Z(x_{t})$
$\frac{1}{\beta}u_{t}(x_{t},y_{t})=\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}+\log Z(x_{t}).$
Since we defined $J_{t}(\pi)$ function,
$\frac{1}{\beta}J_{t}(\pi)=\sum_{y_{t}}\pi(y_{t}|x_{t})\frac{u_{t}(x_{t},y_{t})}{\beta}$
substitute above $\frac{1}{\beta}u_{t}(x_{t},y_{t})$, then we get
$\frac{1}{\beta}J_{t}(\pi)=\sum_{y_{t}}\pi(y_{t}|x_{t})\biggl[\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}+\log Z(x_{t})\biggr].$
@@
To write the above equation in terms of KL-divergence, we decompose the log term as below
@@
Substitute this decomposed term to $\frac{1}{\beta}J_{t}(\pi)$.
$\frac{1}{\beta}J_{t}(\pi)=\sum_{y_{t}}\pi(y_{t}|x_{t})\biggl[\log\frac{\pi(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}-\log\frac{\pi(y_{t}|x_{t})}{\pi^{kl}_{t}(y_{t}|x_{t})}+\log Z(x_{t})\biggr]$
$\frac{1}{\beta}J_{t}(\pi)=D_{KL}(\pi||\pi_{ref,t})-D_{KL}(\pi||\pi^{kl}_{t})+\sum_{y_{t}}\pi(y_{t}|x_{t})\log Z(x_{t}).$
Then, $J_{t}(\pi)$ is redefined as below
$J_{t}(\pi)=\beta\bigl(D_{KL}(\pi||\pi_{ref,t})-D_{KL}(\pi||\pi^{kl}_{t})+\sum_{y_{t}}\pi(y_{t}|x_{t})\log Z(x_{t})\bigr).$
@@
$J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})=\beta\bigl(D_{KL}(\pi^{*}_{t}||\pi_{ref,t})-D_{KL}(\pi^{*}_{t}||\pi^{kl}_{t})-D_{KL}(\pi^{kl}_{t}||\pi_{ref,t})\bigr)$

### v3

@@ B.3 Auxiliary Lemmas
$\hat{Z}(x_{t})=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta_{\mathrm{ref}}}(u_{t}(x_{t},y_{t})-\Delta u_{t}))$
$Z(x_{t})\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta_{\mathrm{ref}}}(\hat{u}_{t}(x_{t},y_{t})-u_{t}(x_{t},y_{t}))),$
$\frac{\hat{Z}(x_{t})}{Z(x_{t})}=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{\hat{u}_{t}(x_{t},y_{t})-u_{t}(x_{t},y_{t})}{\beta_{\mathrm{ref}}}}.$
Therefore,
$\log{\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}}=\log\mathbb{E}_{y_{t}\sim\pi^{kl}(\cdot|x_{t})}[\exp{(-\frac{\Delta u_{t}}{\beta_{\mathrm{ref}}})}]+\frac{1}{\beta_{\mathrm{ref}}}(\Delta u_{t})$
Returning to the definition of KL-divergence,
$D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))=\mathbb{E}_{y_{t}\sim\pi^{kl}}\biggl[\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}\biggr]=\log\mathbb{E}_{y_{t}\sim\pi^{kl}}\biggl[\exp{\bigl(-\frac{\Delta u_{t}}{\beta_{\mathrm{ref}}}\bigr)}\biggr]+\frac{1}{\beta_{\mathrm{ref}}}\mathbb{E}_{\pi^{kl}}[\Delta u_{t}].$
Let $\frac{\Delta u_{t}}{\beta_{\mathrm{ref}}}=X$. Then,
$D_{KL}=-\mathbb{E}[X]+\log\mathbb{E}_{y_{t}\sim\pi^{kl}}\biggl[e^{-(X-\mathbb{E}[X])}\biggr]+\mathbb{E}[X].$
Thus,
$D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))=\log\mathbb{E}_{y_{t}\sim\pi^{kl}}\bigl[e^{-(X-\mathbb{E}[X])}\bigr].$
We now bound $\Delta u_{t}$ in terms of the parameter estimate error. By the definition of $u_{t}$ in (6),
$\Delta u_{t}(y)=\theta^{\top}_{t}\phi(x_{t},y_{t})-\hat{\theta}^{\top}_{t}\phi(x_{t},y_{t})\leq\phi_{max}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}.$
@@
$\frac{\Delta u_{t}}{\beta_{\mathrm{ref}}}\in\biggl[-\frac{\phi_{max}\|\theta_{t}-\hat{\theta}_{t}\|_{2}}{\beta_{\mathrm{ref}}},\frac{\phi_{max}\|\theta_{t}-\hat{\theta}_{t}\|_{2}}{\beta_{\mathrm{ref}}}\biggr].$
Setting $a:=\frac{\phi_{max}\|\theta_{t}-\hat{\theta}_{t}\|_{2}}{\beta_{\mathrm{ref}}}$, we have $X\in[-a,a]$. Since $\mathbb{E}[X]\in[-a,a]$, the centered variable satisfies
@@
Therefore,
$D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))\leq\frac{\phi_{max}^{2}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}^{2}}{2\beta_{\mathrm{ref}}^{2}}.$
B.4 Dynamic regret decomposition
$r_{t}=\{\underbrace{J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})}_{\text{Reference-induced bias}}\}+\{\underbrace{J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})}_{\text{Learning Error}}\}.$ (15)
Learning Error.
$R^{error}_{T}=\sum^{T}_{t=1}\bigl(J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})\bigr).$
By Lemma B.9, with probability at least $1-\delta$,
$R^{error}_{T}\leq\mathcal{O}\bigl(T^{\kappa}V_{T}+T^{1-\frac{\kappa}{2}}\sqrt{\log(T/\delta)}+T^{1-\kappa})$ (16)
where $\kappa\in(0,1)$.
To apply the estimation bound uniformly over all rounds, we invoke Lemma B.9 with confidence level $\delta/T$ and take a union bound over $t=1,\ldots,T$. Therefore, with probability at least $1-\delta$, for all $t$ simultaneously,
$\|\widehat{\theta}_{t}-\theta_{t}\|_{2}\leq\frac{4L_{\sigma}\phi_{\max}^{2}}{m_{0}c}V_{t,W}+\frac{\sqrt{\lambda+4W\phi_{\max}^{2}}}{m_{0}cW}\sqrt{2\left[\frac{d}{2}\log\left(1+\frac{4W\phi_{\max}^{2}}{d\lambda}\right)+\log\left(T/\delta\right)\right]}+\frac{\lambda\theta_{\max}}{m_{0}cW}.$
Proof of Learning Error.
We now establish the bound on $R^{error}_{T}$. We have
$r_{t}=J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})=\sum_{y_{t}\in\mathcal{Y}}\pi^{kl}_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})-\sum_{y_{t}\in\mathcal{Y}}\pi_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t}).$
Thus, we estimate
$\sum_{y_{t}\in\mathcal{Y}}\pi^{kl}_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})-\sum_{y_{t}\in\mathcal{Y}}\pi_{t}(y_{t}|x_{t})\;u_{t}(x_{t},y_{t})\leq\sum_{y_{t}\in\mathcal{Y}}\biggl|\bigl(\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\bigr)\cdot u_{t}(x_{t},y_{t})\biggr|.$
$\sum_{y_{t}\in\mathcal{Y}}\biggl|\bigl(\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\bigr)\cdot u_{t}(x_{t},y_{t})\biggr|\leq\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|\cdot|u_{t}(x_{t},y_{t})|.$
@@ B.4 Dynamic regret decomposition
$\leq\|u_{t}(x_{t},\cdot)\|_{\infty}\sum_{y_{t}\in\mathcal{Y}}\Bigl|\pi_{t}^{kl}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\Bigr|.$
Since $\|u\|_{\infty}\leq\|u\|_{2}\leq\|u\|_{1}$ for any finite vector $u$, it follows that,
$r_{t}\leq\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|\cdot\|u_{t}(x_{t},\cdot)\|_{\infty}\leq\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|\cdot\|u_{t}(x_{t},\cdot)\|_{2}.$
Writing $\sum_{y_{t}\in\mathcal{Y}}|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})|$ as the $L_{1}$ norm $\|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\|_{1}$ and applying Assumption B.3, we obtain the upper bound
$r_{t}\leq\theta_{max}\cdot\phi_{max}\cdot\|\pi^{kl}_{t}(y_{t}|x_{t})-\pi_{t}(y_{t}|x_{t})\|_{1}.$
Applying Pinsker’s inequality to the $L_{1}$ norm and invoking Lemma B.10,
$J_{t}(\pi^{kl}_{t})-J_{t}(\pi_{t})\leq\sqrt{2D_{KL}(\pi^{kl}_{t}(y_{t}|x_{t})||\pi_{t}(y_{t}|x_{t}))}\leq\frac{\theta_{max}\phi^{2}_{max}}{\beta_{\mathrm{ref}}}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}.$
Summing over $t$ and applying Lemma B.9, together with $\sum^{T}_{t=1}V_{t,W}\leq W\;V_{T}$,
$R^{error}_{T}\leq\sum^{T}_{t=1}\frac{\theta_{max}\phi^{2}_{max}}{\beta_{\mathrm{ref}}}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}.$
Using Lemma B.9 uniformly over $t=1,\ldots,T$,
$R_{T}^{\mathrm{error}}\leq\frac{\theta_{\max}\phi_{\max}^{2}}{\beta_{\mathrm{ref}}}\sum_{t=1}^{T}\|\widehat{\theta}_{t}-\theta_{t}\|_{2}.$
Therefore,
$R_{T}^{\mathrm{error}}\leq C\sum_{t=1}^{T}\left[V_{t,W}+W^{-1/2}\sqrt{\log(T/\delta)}+W^{-1}\right],$
where $C$ is a constant that depends on $\theta_{\max},\phi_{\max},L_{\sigma},m_{0},c,\lambda,d,$ and $\beta_{\mathrm{ref}}$. By Lemma B.7,
$\sum_{t=1}^{T}V_{t,W}\leq WV_{T}.$
Hence,
$R_{T}^{\mathrm{error}}\leq C\left[WV_{T}+TW^{-1/2}\sqrt{\log(T/\delta)}+TW^{-1}\right].$
Setting $W=T^{\kappa}$ gives
@@
$R_{T}^{\mathrm{error}}\leq\mathcal{O}\left(T^{\kappa}V_{T}+T^{1-\kappa/2}\sqrt{\log(T/\delta)}+T^{1-\kappa}\right).$
Reference-Induced Bias.
Define the reference-induced bias as
$R^{bias}_{T}=\sum^{T}_{t=1}\bigl(J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})\bigr).$
where $\pi_{t}^{\mathrm{kl}}$ is the idealized KL-regularized comparator. Under Assumption B.5 and the full-support condition on the reference policy, the cumulative reference-induced bias satisfies
$R_{T}^{\mathrm{bias}}\leq\mathcal{O}(1+V_{T}).$
When the initial mismatch term is bounded, the drift-dependent part is $\mathcal{O}(V_{T})$.
Proof of Reference Bias.
Recall the definition of $\pi^{kl}(y_{t}|x_{t})$,
$\pi^{kl}_{t}(y_{t}|x_{t})=\frac{1}{Z(x)}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta_{\mathrm{ref}}}u_{t}(x_{t},y_{t})},$
$\log\pi^{kl}_{t}(y_{t}|x_{t})=\log\pi_{ref,t}(y_{t}|x_{t})+\frac{1}{\beta_{\mathrm{ref}}}u_{t}(x_{t},y_{t})-\log Z(x_{t}).$
Rearranging terms we obtain
$\frac{1}{\beta_{\mathrm{ref}}}u_{t}(x_{t},y_{t})=\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}+\log Z(x_{t}).$
Recalling the definition of $J_{t}(\pi)$ in (9),
$\frac{1}{\beta_{\mathrm{ref}}}J_{t}(\pi)=\sum_{y_{t}}\pi(y_{t}|x_{t})\frac{u_{t}(x_{t},y_{t})}{\beta_{\mathrm{ref}}},$
and substituting the expression for $\frac{1}{\beta_{\mathrm{ref}}}u_{t}(x_{t},y_{t})$ we get
$\frac{1}{\beta_{\mathrm{ref}}}J_{t}(\pi)=\sum_{y_{t}}\pi(y_{t}|x_{t})\biggl[\log\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}+\log Z(x_{t})\biggr].$
@@
To write the above equation in terms of KL-divergence, we decompose the log-ratio as follows:
@@
Substituting into the expression for $\frac{1}{\beta_{\mathrm{ref}}}J_{t}(\pi)$,
$\frac{1}{\beta_{\mathrm{ref}}}J_{t}(\pi)=\sum_{y_{t}}\pi(y_{t}|x_{t})\biggl[\log\frac{\pi(y_{t}|x_{t})}{\pi_{ref,t}(y_{t}|x_{t})}-\log\frac{\pi(y_{t}|x_{t})}{\pi^{kl}_{t}(y_{t}|x_{t})}+\log Z(x_{t})\biggr],$
$\frac{1}{\beta_{\mathrm{ref}}}J_{t}(\pi)=D_{KL}(\pi||\pi_{ref,t})-D_{KL}(\pi||\pi^{kl}_{t})+\sum_{y_{t}}\pi(y_{t}|x_{t})\log Z(x_{t}).$
Therefore,
$J_{t}(\pi)=\beta_{\mathrm{ref}}\bigl(D_{KL}(\pi||\pi_{ref,t})-D_{KL}(\pi||\pi^{kl}_{t})+\sum_{y_{t}}\pi(y_{t}|x_{t})\log Z(x_{t})\bigr).$
@@
$J_{t}(\pi^{*}_{t})-J_{t}(\pi^{kl}_{t})=\beta_{\mathrm{ref}}\bigl(D_{KL}(\pi^{*}_{t}||\pi_{ref,t})-D_{KL}(\pi^{*}_{t}||\pi^{kl}_{t})-D_{KL}(\pi^{kl}_{t}||\pi_{ref,t})\bigr).$